OpenStudy (anonymous):

if limit as x approaches a F(x) + g(x)= 2 and limit as x approaches f(x) - g(x) =1 find the limit f(x)g(x)

OpenStudy (shubhamsrg):

x approaches what ?

OpenStudy (anonymous):

sorry a for both

OpenStudy (shubhamsrg):

f(a) + g(a) = 2 f(a) - g(a) =1 solve for f(a) g(a) .. hope that'd help..

OpenStudy (anonymous):

i dont understand how tho

OpenStudy (anonymous):

3/4

OpenStudy (shubhamsrg):

understand what? how the eqns came? or how to solve the eqns ?

OpenStudy (anonymous):

how to solve or anything can you show steps?

OpenStudy (anonymous):

(f+g)^2 -(f-g)^2 = 4fg -> 2^2-1^1 = 3 => lim fg = 3/4

OpenStudy (anonymous):

but why do you square to f and g ?

OpenStudy (anonymous):

the*